OpenStudy (anonymous):

Is b a linear combination of a1, a2, and a3? a1= (1, -2, 0); a2= (0,1,2); a3=(5, -6, 8); b= (2,-1,6)

OpenStudy (phi):

form a matrix with the a's as vectors, and b as the last vector use elimination on this augmented matrix.

OpenStudy (anonymous):

the last matrix i Have is row 1: 1 0 5 2 row 2: 0 1 4 3 row 3: 0 0 0 0 I don't know what to do from here. The answer says "No, b is not a linear combination of a1, a2, and a3"

OpenStudy (phi):

you have 2 pivot columns. those are your basis the last column says b= 2*1st pivot col + 3*2nd pivot col btw, a3 is a linear combination of a1 and a2

OpenStudy (phi):

You can check 2*a1 + 3*a2 should = b

OpenStudy (phi):

Also, if b were not a combination, it would have been a pivot column

OpenStudy (anonymous):

I see that the back of the book is wrong, but I'm still a little lost as to your explanation. Can you explain "you have 2 pivot...." in a little more detail

OpenStudy (phi):

The pivot is the leading (left most) non-zero entry in a row So the 1 in location (1,1) is a pivot. Also, the 1 in location (2,2). Notice there is only 1 pivot in the 1st column. and one in the 2nd column that means those two vectors are independent

OpenStudy (phi):

the columns that do not have a pivot are dependent. their entries tell the dependence. for example the 2 in location 1,3 tells you that 2 times the first pivot column is part of vector b the 3 in location 2,3 tells you that 3 times the 2nd pivot column is part of b

OpenStudy (phi):

the same with a3. a3= 5*a1 + 4*a2

OpenStudy (anonymous):

location 1,3? isn't that the 5? and location 2,3 is a 4 i believe. Unless I dont understand how the location thing works. The last matrix tells me: x1+ 5x3=2 x2+4x3=3 sorry, I'm still stuck

OpenStudy (phi):

sorry, I can't count. change 3 to 4

OpenStudy (phi):

the 3rd column is a3 and the 4th column is b

OpenStudy (phi):

a1= (1, -2, 0); a2= (0,1,2); a3=(5, -6, 8); b= (2,-1,6) 2*a1= (2, -4, 0) 3*a2= (0, 3, 6) -------------- sum = (2, -1,6) = b

OpenStudy (phi):

btw, normally I would write the vectors vertically, but it's a pain to do so here.

OpenStudy (anonymous):

I guess here is the part that I'm stuck. I see that they can add up to b. But... how I understand it is x1a1+x2a2+x3a3=b so if I solve for x1, x2, and x3. I should have some answer. The two pivot columns tell me that x1 and x2 are basic variables. I still don't see how I can go from my final matrix row 1: 1 0 5 2 row 2: 0 1 4 3 row 3: 0 0 0 0 to solving x1, and x2. I get here x1+ 5x3=2 x2+4x3=3 where x3 is a free variable

OpenStudy (phi):

That is a valid way to think of it. But there is another way, as linear combination of vectors. so the first column (1,0,0) means a1 depends on a1 and 0 times on a2 or a3 the 2nd column (0,1,0) means a2 depends only on a2 (0*a1+1*a2+0*a3) the 3rd column means that a3= 5*a1+4*a2+0*a3 (it is dependent) It's another way to interpret the matrix, but explaining why it works is not so easy for me

OpenStudy (phi):

To actually solve your system, set your free variable to something convenient like zero

OpenStudy (phi):

and to be complete you would add in the null space \[x=\left(\begin{matrix}2 \\ 3 \\ 0\end{matrix}\right)+c\left(\begin{matrix}-5 \\ -4 \\ 1\end{matrix}\right)\]

OpenStudy (phi):

But that's for solving the system. Determining if b is in the column space of A is a simpler matter.